OpenStudy (anonymous):

A triangle has an area of 60in2. Its base is 10 inches and its height is √x inches. What is the value of x? What is the height of the triangle? Please show all work that you use to determine your answer. (Hint: The formula for the area of a triangle is A= 1/2bh where b=base and h=height)

jimthompson5910 (jim_thompson5910):

A= 1/2bh 60 = 1/2*10*sqrt(x) 60*2 = 10*sqrt(x) 10*sqrt(x) = 120 Keep going to solve for x

OpenStudy (anonymous):

the base is 10 in, the height is 12 in @jim_thompson5910

OpenStudy (anonymous):

?*

jimthompson5910 (jim_thompson5910):

the height is 12, so what is x?

OpenStudy (anonymous):

12? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

the height is sqrt(x)

jimthompson5910 (jim_thompson5910):

you found the height to be 12, so the two are equal, ie sqrt(x) = 12 so x must be what?

OpenStudy (anonymous):

10? ahaha i suck at everything .. :/

jimthompson5910 (jim_thompson5910):

don't say that